From GlueXWiki
Jump to: navigation, search

April 9, 2015 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. Gas puzzle
    • CO2 monitoring [2]
  2. CDC update (Beni, Mike)
  3. FDC update (Lubomir)
    • Alignment status [3]
  4. Electronics (Fernando, Cody, Chris, Nick)
  5. Engineering
  6. Other


Participants: Naomi(CMU), Sean(NU), Fernando, Luke, Mike, Dave, Chris, Nick, Cody, Beni, Simon, and Lubomir (JLab).

Gas puzzle

- The CO2 sensor was moved last Friday to the Hall connected to the CDC (figure attached above). Beni changed the CO2 gas factor from 0.74 to 1 on Sunday: the CO2 percentage started going down. On Tuesday Beni changed the gas factor to 0.5 (just a guess) then CO2 started going up in the right direction.

- Beni did several other tests: Ar gas factor set from 1.391 to 1 for a short period of time. Not clear yet what the problem is with the gas mixture.

CDC update

- There are many files collected with the BCAL cosmics trigger to be used for the CDC alignment. They are being analyzed now.

FDC update

- Lubomir finished aligning the individual packages - there are 48 parameters per package, 8 per cell. Next step: use three packages for tracking and aligning the fourth as one piece. Using the outer packages for tracking, there's an improvement of the resolution compared to the resolution obtained by using the package itself (figure attached). The new resolution is very close to the one obtained in the EEL where the timing from the scintillators that are very close to the chamber, is the best.


- The HV filter boxes have arrived.

- Nick and Chris will help with the spare package cabling and installation of the fADCs, HV and LV modules, as instructed by Fernando.


- The FDC spare package was moved to the Tagger successfully (picture attached). It will be used to test it as a transition radiation detector.